Brief Exercise 6-12 On June 30, Calico Fabrics has the following data pertaining to the retail inventory method. Goods available for sale: at cost $38,000; at retail $50,000; net sales $40,000; and ending inventory at retail $10,000.

Compute the estimated cost of the ending inventory using the retail inventory method.

The estimated cost of the ending inventory $______.
Answer #1

Cost to retail percentage = cost/retail price = 38000/50000 = 0.76

cost of sales = net sales*cost to retail percentage = 40000*0.76 = 30400

ending inventory = cost of goods available for sale - cost of sales = 38000-30400 = 7600